# On Long-Term Problems in Multiplicative Ideal Theory and Factorization Theory

### v2

@@ 2. Preliminaries
Let $\mathbb{N}$ denote the set of positive integers and set $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For integers $a,b\in\mathbb{Z}$, let $[a,b]=\{x\in\mathbb{Z}\colon a\leq x\leq b\}$ denote the discrete interval between $a$ and $b$. For subsets $A,B$ of an additive abelian group $G$, let $A+B=\{a+b\colon a\in A,b\in B\}\subseteq G$ denote their sumset. If $S$ is any set, then $\mathcal{P}(S)$ denotes the system of all subsets of $S$ and $\mathcal{P}_{\text{\rm fin}}(S)$ the system of all finite, nonempty subsets of $S$.
@@
Let $r$ be an ideal system on $M$. A subset $I\subseteq M$ is called an $r$-ideal if $I_{r}=I$. We denote by $\mathcal{I}_{r}(M)$ the set of all nonempty $r$-ideals, and we define $r$-multiplication by setting $I\cdot_{r}J=(IJ)_{r}$ for all $I,J\in\mathcal{I}_{r}(M)$. Then $\mathcal{I}_{r}(M)$ together with $r$-multiplication is a reduced semigroup with identity element $M$.
@@
All rings are commutative and have an identity element. For a commutative ring $D$, we denote by $D^{\times}$ its group of units and by $D^{\bullet}$ its monoid of regular elements. Let $D$ be a domain with quotient field $K$. We denote by $\mathcal{A}(D)$ the set of atoms, by $\mathsf{q}(D)$ the quotient field of $D$, by $\overline{D}$ the integral closure of $D$, and by $\widehat{D}=\widehat{D^{\bullet}}\cup\{0\}$ the complete integral closure of $D$.
@@ 3. Intersecting Valuation Domains
Let $V$ be a valuation domain. For a polynomial $f\in V[X]$, let $c(f)$ be the ideal of $V$ generated by the coefficients of $f$. Let $N=\{f\in V[X]\mid c(f)=V\}$ and finally, let $V(X)=N^{-1}V[X]$. $V(X)$ is known as the Nagata extension of $V$. $V(X)$ is a valuation domain which is almost identical to $V$. The only significant difference between $V$ and $V(X)$ is that $V(X)$ has more units.
@@ 4. Classification of Integral Domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$
We survey recent developments in the classification of domains that lie strictly between the polynomial ring $\mathbb{Z}[X]$ and its field of fractions $\mathbb{Q}[X]$. We focus on two major studies:
@@
  • ¯$\overline{\mathbb{Z}}_{p}$ denotes the absolute integral closure of $\mathbb{Z}_{p}$.
  • ^¯$\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}}_{p}$.
@@
Let $\underline{E}=\prod_{p}E_{p}$ be a subset of $\overline{\widehat{\mathbb{Z}}}$. Following [117, Definition 2.11], we say that $\underline{E}$ is polynomially factorizable if, for every $g\in\mathbb{Z}[X]$ and every $\alpha=(\alpha_{p})\in\underline{E}$, there exist $n,d\in\mathbb{Z}$ with $n,d\geq 1$ such that $g(\alpha)^{n}/d$ is a unit of $\overline{\widehat{\mathbb{Z}}}$.
@@
This result completes the classification of Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ and provides a powerful characterization of Dedekind domains in this setting, linking local valuation data and integer-valued polynomial constraints to global structural properties such as the class group. The result is obtained by characterizing discrete valuation rings (DVRs) $W$ of the field $\mathbb{Q}(X)$ that are residually algebraic over $\mathbb{Z}_{(p)}$ for some prime $p$ (see [118, Corollary 2.28]).
@@
  • Krull or ((((strong)))) Mori domains: Extending classification results to settings with additional constraints (e.g., finite residue fields, specific $v$-class groups, etc.).
@@ 5. Generalizing Kronecker Function Domains
  • One could wonder whether the base ring $D$ is essential. Is it possible to build a Kronecker function domain using just $K$ without $D$?
@@
Halter-Koch’s notion of Kronecker domain has been known since 2003 and very little has been done in that time to study them. Actually, Halter-Koch himself extended what is known by himself by demonstrating that they can be built with a variant of the conventional construction ([88, Theorem 6.5.5]). Many questions could be posed. We will content ourselves with one.
@@ 7. Theory of $w$-modules
    An $R$-module $M$ is called GV-torsion-free if, for any $J\in\operatorname{GV}(R)$, the condition $Jx=0$ (for some $x\in M$) implies $x=0$. On the other hand, $M$ is called GV-torsion if, for every nonzero element $x\in M$, there exists $J\in\operatorname{GV}(R)$ such that $Jx=0$. It is evident that the pair (GV-torsion modules, GV-torsion-free modules) defines a hereditary torsion theory, first introduced in [102, p. 1618].
@@
It is well-known that Noetherian rings play an important role in commutative algebra and algebraic geometry. Using $w$-operations, strong Mori domains were introduced in [140]. It was previously claimed ([57, Corollary IX.5.5]) that for a domain $D$ with quotient field $K$, the injective module $E(K/D)$ is $\Sigma$-injective if and only if $D$ is a Mori domain. However, this statement is incorrect: one must replace ‘Mori’ with ‘strong Mori’ (see [56, Theorem 9]). In recent years, many classical theorems about Noetherian rings have been generalized to strong Mori domains.
@@
Analogous to how Dedekind domains can be characterized by the invertibility of nonzero ideals, Krull domains can be characterized as follows ([101, Theorem 3.6], [139, Theorem 7.9.3]):
@@
    The $w$-factor ring $(D/I)_{w}$ is a zero-dimensional semilocal principal ideal ring for every proper $w$-ideal $I$ of $D$, [26, Theorem 4.5].
@@
    $(D/I)_{w}$ is a quasi-Frobenius ring [148, Theorem 9].
[…]
@@
[…]
@@ 9. Gorenstein Multiplicative Ideal Theory
    such that $M\cong\ker(F^{0}\to F^{1})$ and exactness is preserved under the functor $-\otimes_{R}E$ for every injective $R$-module $E$.
@@
For example, in the classical setting, a domain $D$ is Prüfer if and only if it is locally a valuation domain ([139, Corollary 3.7.22]). In the Gorenstein setting, while every G-Prüfer domain is locally G-Prüfer [145, Corollary 11], the converse fails in general [147, Example 4.1]. Hence, the direct classical analogy does not hold perfectly in Gorenstein contexts.
Classically, $D$ is Prüfer if and only if every overring of $D$ is flat [107, Theorem 6.10]. Analogously, it has been shown that if $R$ is G-Prüfer, then every overring of $R$ is Gorenstein flat [145, Theorem 6]. However, it remains open whether the converse also holds. Namely, is a domain $D$ G-Prüfer if every overring of $D$ is Gorenstein flat? This is a natural Gorenstein-homological parallel to a well-known Prüfer criterion, but it awaits resolution.
A key feature of Dedekind domains $D$ is that every nonzero ideal is invertible, which is equivalent to demanding that $D/I$ be quasi-Frobenius for each nonzero ideal $I$ ([148, Theorem 13]). Similarly, in the Gorenstein Dedekind setting, one has:
  • An integral domain $D$ is G-Dedekind if and only if, for every nonzero element $a\in D$, the factor ring $D/(a)$ is quasi-Frobenius ([139, Theorem 11.7.7]). This implies that every nonzero ideal of $D$ is a $v$-ideal. In fact, $D$ is a G-Dedekind domain if and only if it is a Noetherian divisorial domain.
  • In general, a G-Dedekind domain is one-dimensional and Noetherian ([139, Theorem 11.7.7 and Corollary 11.7.8]), yet can exhibit different behaviors regarding integral closure, reminiscent of how Gorenstein rings relax some classical regularity or integrality conditions.
Recent connections show that factor rings of the form $(D/xD)_{w}$ being quasi-Frobenius for all nonzero $x$ in $D$ characterizes G-Krull domains [146, Theorem 2.7]. Such Gorenstein-enabled statements suggest deep analogies between multiplicative ideal theory and Gorenstein homological algebra, but also underscore certain exceptions where classical and Gorenstein theories diverge.
@@ 10. Cotorsion Theory
    =$\mathcal{B}\;=\;\bigl{\{}\,N\;\bigm{|}\;\operatorname{Ext}^{1}_{R}(A,N)=0\text{ for all }A\in\mathcal{A}\bigr{\}}.$
@@
    Over a principal ideal domain $D$, the cotorsion pair $(\mathcal{A},\mathcal{B})=(\mathrm{Flat},\mathrm{Cotorsion})$ satisfies:
@@
A variant of cotorsion theory, called Tor-cotorsion theory, shifts attention from $\operatorname{Ext}$ to $\operatorname{Tor}$ and hence emphasizes torsion properties and tensor products.
@@
An $R$-module $M$ is called Tor-cotorsion if
$\operatorname{Tor}_{1}^{R}(F,M)\;=\;0\quad\text{for every flat $R$-module }F.$
In classical cotorsion theory, one requires $\operatorname{Ext}^{1}_{R}(F,M)=0$ for every flat $R$-module $F$. In Tor-cotorsion theory, one instead demands that $\operatorname{Tor}_{1}^{R}(F,M)=0$ for every flat $R$-module $F$. Both conditions provide different perspectives on how flat modules (or certain torsion-free classes) might interact with a given module, illuminating complementary aspects of homological dimension and torsion phenomena.
@@
Let $\mathscr{F}$ be the class of all flat $R$-modules. Then $(\mathscr{F},\mathcal{M})$ is a (classical) Tor-cotorsion pair in the category of $R$-modules. In a Tor-cotorsion context, we focus on $R$-modules $M$ with $\operatorname{Tor}_{1}^{R}(F,M)=0$ rather than on $\operatorname{Ext}^{1}_{R}(F,M)=0$.
Cotorsion theory (and its Tor-cotorsion counterpart) can effectively describe structural aspects of rings that are central in multiplicative ideal theory. For instance:
@@
    An integral domain $D$ is Krull if and only if the global weak $w$-projective dimension $w.w.gl.\dim(D)$ is at most $1$; see [141, Theorem 4.3].
@@
[…]
@@
Develop explicit cotorsion and Tor-cotorsion characterizations of rings and integral domains that arise in multiplicative ideal theory. In particular, investigate how homological invariants such as
[…]
@@
[…]
@@ 11. Factorization Theory in Mori Monoids
We start with a lemma highlighting the abundance of Mori and Krull monoids in commutative ring theory. For more on the ideal theory of Mori monoids we refer to [88, Chapter 5] and for more of this flavor in the setting of $v$-Marot and $t$-Marot rings see [27, 14]. For Mori monoids with a combinatorial flavor see [42]. Moreover, we refer to a survey of Facchini [43] for the role of Krull monoids in module theory.
skip
@@
    Let $M$ be a Mori monoid. If $M$ is seminormal or if the conductor $(M\colon\widehat{M})$ is nonempty, then the complete integral closure $\widehat{M}$ is Krull. Furthermore, the complete integral closure of a strong Mori domain is Krull.
@@
For (1) and (2) we refer to [70] and for (3) to [69, Theorem 4.5]. The first claim in (4) is the Theorem of Mori-Nagata and the second claim follows from [65, Example 2.1].
(5) For a seminormal $M$ the claim follows from [88, Theorem 5.2.8] and if the conductor is nonempty, then it follows from [64, Theorem 2.3.5]. If $M$ is a strong Mori domain, then the claim follows from [139, Theorems 7.7.14 and 7.10.21]. ∎
@@
    All length sets are finite and nonempty (((([64, Theorem 2.2.9])))).
@@
    $\omega(a)<\infty$ for all $a\in M$ (((([65, Theorem 4.2])))).
@@
$M\ \text{is half-factorial}\quad\text{if and only if}\quad\Delta(M)=\emptyset\quad\text{if and only if}\quad\rho(M)=1\,,$
@@
Consider two simple examples of Mori domains, Cohen-Kaplansky domains and one-dimensional local Mori domains. Their arithmetic is extremely simple and it stimulates research to which extent the arithmetic of more general domains fulfills some resp. all of their arithmetic finiteness properties. Let $D$ be a domain and let $r$ be an ideal system on $D$ such that $\mathcal{I}_{r}(D)\subseteq\mathcal{I}(D)$. Then $D$ is called a Cohen-Kaplansky domain if it satisfies one of the following equivalent conditions ([71, Proposition 4.5])
@@
(1) The statement on the elasticity follows from [2, Theorem 2.12] and from [64, 2.9.2+2.10.7]. The other statements follow from [61, Theorem 4.1] and from [72, Theorem 3.9].
(2) This follows from [64, 3.1.4+4.4.11] and from [138]. ∎

### v3

# arXiv:2502.21020v3[math.AC] 28 Feb 2026

@@ 2. Preliminaries
Let $\mathbb{N}$ denote the set of positive integers and set $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For integers $a,b\in\mathbb{Z}$, let $a,b]=\{x\in\mathbb{Z}\colon a\leq x\leq b\}$ denote the discrete interval between $a$ and $b$. For subsets $A,B$ of an additive abelian group $G$, let $A+B=\{a+b\colon a\in A,b\in B\}\subseteq G$ denote their sumset. If $S$ is any set, then $\mathcal{P}(S)$ denotes the system of all subsets of $S$ and $\mathcal{P}_{\text{\rm fin}}(S)$ the system of all finite, nonempty subsets of $S$.
@@
Let $r$ be an ideal system on $M$. A subset $I\subseteq M$ is called an $r$-ideal if $I_{r}=I$. We denote by $\mathcal{I}_{r}(M)$ the set of all nonempty $r$-ideals, and we define $r$-multiplication by setting $I\cdot_{r}J=(IJ)_{r}$ for all $I,J\in\mathcal{I}_{r}(M)$. Then $\mathcal{I}_{r}(M)$ together with $r$-multiplication is a reduced semigroup with identity element $M$, and it is a multiplicative lattice. Conversely, by [39], for every multiplicative lattice there are a commutative semigroup $H$ and a weak ideal system $r$ on $H$ such that $L$ is isomorphic to the lattice of $r$-ideals of $H$.
@@
[…]
@@ 3. Intersecting Valuation Domains
Let $V$ be a valuation domain. For a polynomial $f\in V[X]$, let $c(f)$ be the ideal of $V$ generated by the coefficients of $f$. Let $N=\{f\in V[X]\mid c(f)=V\}$ and finally, let $N^{-1}V[X]. V(X)V(X)$ is known as the Nagata extension of $V$. $V(X)$ is a valuation domain which is almost identical to $V$. The only significant difference between $V$ and $V(X)$ is that $V(X)$ has more units.
@@ 4. Classification of Integral Domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$
We survey recent developments in the classification of domains that lie strictly between the polynomial ring $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$. We focus on two major studies:
@@
  • $\overline{\mathbb{Z}}_{p}$ denotes the absolute integral closure of $\mathbb{Z}_{p}$.
  • $\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}}_{p}$.
@@
Let $\underline{E}=\prod_{p}E_{p}$ be a subset of $\overline{\widehat{\mathbb{Z}}}$. Following [118, Definition 2.11], we say that $\underline{E}$ is polynomially factorizable if, for every $g\in\mathbb{Z}[X]$ and every $\alpha=(\alpha_{p})\in\underline{E}$, there exist $n,d\in\mathbb{Z}$ with $n,d\geq 1$ such that $g(\alpha)^{n}/d$ is a unit of $\overline{\widehat{\mathbb{Z}}}$.
@@
This result completes the classification of Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ and provides a powerful characterization of Dedekind domains in this setting, linking local valuation data and integer-valued polynomial constraints to global structural properties such as the class group. The result is obtained by characterizing discrete valuation rings (DVRs) $W$ of the field $\mathbb{Q}(X)$ that are residually algebraic over $\mathbb{Z}_{(p)}$ for some prime $p$ (see [119, Corollary 2.28]).
@@
  • Krull or ((strong)) Mori domains: Extending classification results to settings with additional constraints (e.g., finite residue fields, specific $v$-class groups, etc.).
@@ 5. Generalizing Kronecker Function Domains
  • One could wonder whether the base ring $D$ is essential. Is it possible to build a Kronecker function domain using just $K$ - without $D$?
@@
Halter-Koch’s notion of Kronecker domain has been known since 2003 and very little has been done in that time to study them. Actually, Halter-Koch himself extended what is known by himself by demonstrating that they can be built with a variant of the conventional construction ([90, Theorem 6.5.5]). Many questions could be posed. We will content ourselves with one.
@@ 7. Theory of $w$-modules
    An $R$-module $M$ is called GV-torsion-free if, for any $J\in\operatorname{GV}(R)$, the condition $Jx=0$ (for some $x\in M$) implies $x=0$. On the other hand, $M$ is called GV-torsion if, for every nonzero element $x\in M$, there exists $J\in\operatorname{GV}(R)$ such that $Jx=0$. It is evident that the pair (GV-torsion modules, GV-torsion-free modules) defines a hereditary torsion theory, first introduced in [104, p. 1618].
@@
It is well-known that Noetherian rings play an important role in commutative algebra and algebraic geometry. Using $w$-operations, strong Mori domains were introduced in [141]. It was previously claimed ([59, Corollary IX.5.5]) that for a domain $D$ with quotient field $K$, the injective module $E(K/D)$ is $\Sigma$-injective if and only if $D$ is a Mori domain. However, this statement is incorrect: one must replace ‘Mori’ with ‘strong Mori’ (see [58, Theorem 9]). In recent years, many classical theorems about Noetherian rings have been generalized to strong Mori domains.
@@
Analogous to how Dedekind domains can be characterized by the invertibility of nonzero ideals, Krull domains can be characterized as follows ([103, Theorem 3.6], [140, Theorem 7.9.3]):
@@
    The $w$-factor ring $(D/I)_{w}$ is a zero-dimensional semilocal principal ideal ring for every proper $w$-ideal II of $D$, [26, Theorem 4.5].
@@
    $(D/I)_{w}$ is a quasi-Frobenius ring [149, Theorem 9].
These results highlight a strong parallel between Dedekind and Krull domains when analyzed through $w$-operations and $w$-ideal structures. This raises the question of whether PvvMDs can be characterized by their $w$-factor rings, where an integral domain $D$ is called a Prüfer $v$-multiplication domain (PvvMD for short) if every nonzero finitely generated ideal of $D$ is $t$-invertible (or $w$-invertible).
@@
Investigate whether the PvvMDs $D$ are precisely those domains for which their $w$-factor rings $(D/I)_{w}$ are FC-rings for every proper $w$-ideal II of $D$, where a commutative ring $R$ is an FC-ring if it is coherent and self-FP-injective.
@@ 9. Gorenstein Multiplicative Ideal Theory
    such that $M\cong\ker(F^{0}\to F^{1})$ and exactness is preserved under the functor $\otimes_{R}E$ for every injective $R$-module $E$.
@@
For example, in the classical setting, a domain $D$ is Prüfer if and only if it is locally a valuation domain ([140, Corollary 3.7.22]). In the Gorenstein setting, while every G-Prüfer domain is locally G-Prüfer [146, Corollary 11], the converse fails in general [148, Example 4.1]. Hence, the direct classical analogy does not hold perfectly in Gorenstein contexts.
Classically, $D$ is Prüfer if and only if every overring of $D$ is flat [109, Theorem 6.10]. Analogously, it has been shown that if $R$ is G-Prüfer, then every overring of $R$ is Gorenstein flat [146, Theorem 6]. However, it remains open whether the converse also holds. Namely, is a domain $D$ G-Prüfer if every overring of $D$ is Gorenstein flat? This is a natural Gorenstein-homological parallel to a well-known Prüfer criterion, but it awaits resolution.
A key feature of Dedekind domains $D$ is that every nonzero ideal is invertible, which is equivalent to demanding that $D/I$ be quasi-Frobenius for each nonzero ideal $I$ ([149, Theorem 13]). Similarly, in the Gorenstein Dedekind setting, one has:
  • An integral domain $D$ is G-Dedekind if and only if, for every nonzero element $a\in D$, the factor ring $D/(a)$ is quasi-Frobenius ([140, Theorem 11.7.7]). This implies that every nonzero ideal of $D$ is a $v$-ideal. In fact, $D$ is a G-Dedekind domain if and only if it is a Noetherian divisorial domain.
  • In general, a G-Dedekind domain is one-dimensional and Noetherian ([140, Theorem 11.7.7 and Corollary 11.7.8]), yet can exhibit different behaviors regarding integral closure, reminiscent of how Gorenstein rings relax some classical regularity or integrality conditions.
Recent connections show that factor rings of the form $(D/xD)_{w}$ being quasi-Frobenius for all nonzero $x$ in $D$ characterizes G-Krull domains [147, Theorem 2.7]. Such Gorenstein-enabled statements suggest deep analogies between multiplicative ideal theory and Gorenstein homological algebra, but also underscore certain exceptions where classical and Gorenstein theories diverge.
@@ 10. Cotorsion Theory
    $\mathcal{B}\;=\;\bigl\{\,N\;\bigm|\;\operatorname{Ext}^{1}_{R}(A,N)=0\text{ for all }A\in\mathcal{A}\bigr\}.$
@@
    Over any ring $D$, the cotorsion pair $(\mathcal{A},\mathcal{B})=(\mathrm{Flat},\mathrm{Cotorsion})$ satisfies:
@@
A variant of cotorsion theory, called Tor-torsion theory, shifts attention from $\operatorname{Ext}$ to $\operatorname{Tor}$ and hence emphasizes torsion properties and tensor products.
@@
Let $\mathcal{A}$ and $\mathcal{B}$ be two classes of modules. The pair $(\mathcal{A},\mathcal{B})$ is called a Tor-torsion pair (or Tor-torsion theory) if
$\mathcal{A}=\mathcal{B}^{\top}\quad\text{and}\quad\mathcal{B}=\mathcal{A}^{\top},$
where
$\mathcal{A}^{\top}=\left\{B\in\mathcal{M}\middle|\operatorname{Tor}_{1}^{R}(A,B)=0\text{ for every }A\in\mathcal{A}\right\}.$
In classical cotorsion theory, one requires $\operatorname{Ext}^{1}_{R}(F,M)=0$ for every flat $R$-module $F$. In Tor-torsion theory, one instead demands that $\operatorname{Tor}_{1}^{R}(F,M)=0$ for every flat $R$-module $F$. Both conditions provide different perspectives on how flat modules (or certain torsion-free classes) might interact with a given module, illuminating complementary aspects of homological dimension and torsion phenomena.
@@
Let $\mathscr{F}$ be the class of all flat $R$-modules. Then $(\mathscr{F},\mathcal{M})$ is a (classical) Tor-torsion pair in the category of $R$-modules. In a Tor-torsion context, we focus on $R$-modules $M$ with $\operatorname{Tor}_{1}^{R}(F,M)=0$ rather than on $\operatorname{Ext}^{1}_{R}(F,M)=0$.
Cotorsion theory (and its Tor-torsion counterpart) can effectively describe structural aspects of rings that are central in multiplicative ideal theory. For instance:
@@
    An integral domain $D$ is Krull if and only if the global weak $w$-projective dimension $w.w.gl.\dim(D)$ is at most $1$; see [142, Theorem 4.3].
@@
The interplay between cotorsion theory and $w$-theory highlights how star operations can refine homological statements, particularly for non-Noetherian rings or those with specialized divisorial properties. For example, focusing on $w$-flatness and $w$-cotorsion can unify certain results in multiplicative ideal theory with relative homological vanishing conditions. In this direction, the cases of PvvMDs and Krull domains have been studied in [122, 123].
@@
Develop explicit cotorsion and Tor-torsion characterizations of rings and integral domains that arise in multiplicative ideal theory. In particular, investigate how homological invariants such as
$\mathrm{Ext}^{1}_{R}(-,-),\quad\mathrm{Tor}_{1}^{R}(-,-),\quad\text{global dimensions, etc.}$
@@
  • Other star-operation domains: Characterize specialized classes (e.g., Gorenstein domains, G-Prüfer, PvvMDs, etc.) via novel cotorsion or Tor-torsion properties.
@@ 11. Factorization Theory in Mori Monoids
We start with a lemma highlighting the abundance of Mori and Krull monoids in commutative ring theory. For more on the ideal theory of Mori monoids we refer to [90, Chapter 5] and for more of this flavor in the setting of $v$-Marot and $t$-Marot rings see [27, 14]. For Mori monoids with a combinatorial flavor see [44]. Moreover, we refer to a survey of Facchini [45] for the role of Krull monoids in module theory.
@@
    Let $M$ be a Mori monoid. If $M$ is seminormal or if the conductor (M:M^)($\colon\widehat{M}$) is nonempty, then the complete integral closure $\widehat{M}$ is Krull. Furthermore, the complete integral closure of a strong Mori domain is Krull.
@@
For (1) and (2) we refer to [72] and for (3) to [71, Theorem 4.5]. The first claim in (4) is the Theorem of Mori-Nagata and the second claim follows from [67, Example 2.1].
(5) For a seminormal $M$ the claim follows from [90, Theorem 5.2.8] and if the conductor is nonempty, then it follows from [66, Theorem 2.3.5]. If $M$ is a strong Mori domain, then the claim follows from [140, Theorems 7.7.14 and 7.10.21]. ∎
@@
    All length sets are finite and nonempty (([66, Theorem 2.2.9])).
@@
    $\omega(a)<\infty$ for all $a\in M$ (([67, Theorem 4.2])).
@@
[…]
@@
Consider two simple examples of Mori domains, Cohen-Kaplansky domains and one-dimensional local Mori domains. Their arithmetic is extremely simple and it stimulates research to which extent the arithmetic of more general domains fulfills some resp. all of their arithmetic finiteness properties. Let $D$ be a domain and let $r$ be an ideal system on $D$ such that $\mathcal{I}_{r}(D)\subseteq\mathcal{I}(D)$. Then $D$ is called a Cohen-Kaplansky domain if it satisfies one of the following equivalent conditions ([73, Proposition 4.5])
@@
(1) The statement on the elasticity follows from [2, Theorem 2.12] and from [66, 2.9.2+2.10.7]. The other statements follow from [63, Theorem 4.1] and from [74, Theorem 3.9].
(2) This follows from [66, 3.1.4+4.4.11] and from [139]. ∎